OpenStudy (laddiusmaximus):

find the center and radius of the given circle. x^2 +y^2+6x-8y-39=0

OpenStudy (mertsj):

Write it like this? (x^2+6x + )+(y^2-8y + )=39

OpenStudy (mertsj):

Now complete the square for the x part and again or the y part. Don't forget to add the numbers to both sides!

OpenStudy (laddiusmaximus):

@Mertsj how do I complete the square?

OpenStudy (mertsj):

Divide the coefficient of x by 2. What do you get?

OpenStudy (laddiusmaximus):

for x^2-6x it would be 3 and for y^2-8y it would be 4 right?

OpenStudy (mertsj):

Now square the 3 and the 4 and add those numbers to both sides.

OpenStudy (mertsj):

(x^2+6x+9)+(y^2-8y+16)=39+9+16

OpenStudy (mertsj):

Now factor the two trinomials.

OpenStudy (laddiusmaximus):

why did I divide by two? common factor?

OpenStudy (mertsj):

Because the relationship between the coefficient of x and the constant term in a trinomial square is always: ( 1/2 the coefficient of x )^2= the constant term.

OpenStudy (mertsj):

\[x^2-6x+9=(x-3)^2 and (\frac{6}{2})^2=9\]

OpenStudy (laddiusmaximus):

ok so I have (x-3)^2+(x+4)^2 = 8^2

OpenStudy (mertsj):

\[x^2+10x+25=(x+5)^2 and (\frac{10}{2})^2=25\]

OpenStudy (mertsj):

yep

OpenStudy (mertsj):

Oh wait. I think it is( y-4)^2

OpenStudy (mertsj):

\[(x+3)^2+(y-4)^2=64\]

OpenStudy (mertsj):

So the center is????

OpenStudy (laddiusmaximus):

(-3,4) r=8

OpenStudy (mertsj):

yep

OpenStudy (mertsj):

Good job!!

OpenStudy (laddiusmaximus):

thanks. I like to be walked though these when I don't know them well.

OpenStudy (mertsj):

yw

OpenStudy (laddiusmaximus):

wait why was it x-4^2?

OpenStudy (mertsj):

It is (y-4)^2 because the factors of y^2-8y+16 are (y-4)^2

OpenStudy (mertsj):

0 (x^2+6x+9)+(y^2-8y+16)=39+9+16

OpenStudy (laddiusmaximus):

ok. gotcha

OpenStudy (mertsj):

Good.

OpenStudy (laddiusmaximus):

@Mertsj can you walk me through one more to make sure I understand the concept?

OpenStudy (mertsj):

Yep

OpenStudy (mertsj):

If you post it.